OpenStudy (anonymous):

Prove: tan2 θ•cos2 θ + cos2 θ = 1. You must show all work.

OpenStudy (solomonzelman):

tan²x=sin²x / cos²x Re-write tan²x in terms of sines and cosine,s cancel the cos²x out, and you get sin²x+cos²x=1

OpenStudy (anonymous):

@SolomonZelman s cancel?

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \tan^2θ\cos^2θ+\cos^2θ=1}\) \(\normalsize\color{blue}{ \frac{\Large{sin^2θ}}{\Large{cos^2θ}}cos^2θ+\cos^2θ=1}\)

OpenStudy (solomonzelman):

cos²θ cancel, and you get an identity, ` sin²x+cos²x=1 `

OpenStudy (anonymous):

how does that prove it?

OpenStudy (solomonzelman):

sin²x+cos²x=1 is an identity, and once you get to that, you can say it is proved

OpenStudy (anonymous):

can you help with another?

OpenStudy (solomonzelman):

I can try :)

OpenStudy (anonymous):

i asked in another question